Job Costs Using Activity-Based Costing

Heitger Company is a job-order costing firm that uses activity-based costing to apply overhead to jobs. Heitger identified three overhead activities and related drivers. Budgeted information for the year is as follows:

Activity Cost Driver Amount of Driver
Materials handling $56,250 Number of moves 2,500
Engineering 142,500 Number of change orders 9,500
Other overhead 209,100 Direct labor hours 51,000

Heitger worked on four jobs in July. Data are as follows:

Job 13-43 Job 13-44 Job 13-45 Job 13-46
Beginning balance $21,400 $19,400 $4,500 $0
Direct materials $5,600 $9,700 $12,700 $9,600
Direct labor cost $810 $990 $1,700 $110
Job 13-43 Job 13-44 Job 13-45 Job 13-46
Number of moves 43 48 27 7
Number of change orders 32 44 24 15
Direct labor hours 810 990 1,700 110

By July 31, Jobs 13-43 and 13-44 were completed and sold. Jobs 13-45 and 13-46 were still in process.

Required:

1. Calculate the activity rates for each of the three overhead activities. Round all activity rates to the nearest cent.

Materials handling rate $ per move
Engineering rate $ per change order
Other overhead rate $ per direct labor hour

1. In job-order costing, departmental overhead rates and activity-based costing affect only the application of overhead and is put on the job-order costing sheet and must include all drivers for which overhead is applied.

2. Prepare job-order cost sheets for each job showing all costs through July 31.

When required, round your answers to the nearest dollar. If an amount is zero, enter "0".

Heitger Company
Job-Order Cost Sheets
Job 13-43 Job 13-44 Job 13-45 Job 13-46
Balance, July 1 $ $ $ $
Direct materials
Direct labor cost
Materials handling
Engineering
Other overhead
Total $ $ $ $

2. The job-order cost sheet must include a unique number or name for this particular job, all direct costs, and all overhead costs associated with the job.

3. Calculate the balance in Work in Process on July 31.
$____

4. Calculate the cost of goods sold for July.
$____

5. What if Job 13-46 required no engineering change orders? What is the difference in the new cost of Job 13-46?

$____ lower

How would the cost of the other jobs be affected?

no effect
Answer #1

The answer has been presented in the supporting sheet. All the parts has been solved with detailed explanation and calculation. For detailed answer refer to the supporting sheet.

Answer e Part 1) Activity 5 Material Handling 5 Engineering - Other Overhead Cost Per Activity Cost 56250 142500 209100 Amoun

26 Part 3) 27 Balance in work in process 28 = Job 13-45+ Cost of Job 1346 29 = 26837.5 +10543.5 30 = $ 37381 32 Part 4) 33 Th